Show how-3.75 is represented in a computer architecture that uses a floating point architecture with 7 bits exponent and 8 bits fraction and 1 bit sign
Answer #1

Answer: 1 0100000 11100000

Explanation:

-3.75

3.75 = 11. 11 = 1.111 * 21

Sign bit = 1; since negative

Mantessa = 11100000

Exponent = bias + power of 2 (bias of 16 bit floating point representation is 63

Therefore Exponent = 63+1 = 64 = 0100000

The 16 bit floating point number for -3.75 is 1 0100000 11100000